OpenStudy (anonymous):

solve (2^5/2)-(2^3/2)

OpenStudy (anonymous):

not an available answer, either it is 2 or 2 with a fractional exponent, I have forgotten how to deal with addition and subtraction of fractional exponents...

OpenStudy (anonymous):

how do you reduce the fraction 5/2 to 4 and the fraction 3/2 to 2 ?

OpenStudy (anonymous):

(2^5)/2)-(2^3)/2) 16 - 4 12 ---------------- (2^(5/2)) - (2^(3/2)) ( 4 Sqrt[2]) - ( 2 Sqrt[2] ) 2 Sqrt[2]

OpenStudy (anonymous):

not an available answer... it is 2 or some power of 2...the question is: 2 raised to the power 5/2 minus 2 to the power 3/2

OpenStudy (anonymous):

\[2^{5/2}-2^{3/2}=2^{3/2} \]

OpenStudy (anonymous):

That is the answer, but how did you get that ?

OpenStudy (anonymous):

see follow the rule BODMAS i.e. bracket,order,division,multiplication,addition,subtraction now here use the exponential identity i.e. a^m - a^n = a^(m-n) =(2^5/2) - (2^3/2) = 2^5/2 - 2^3/2 = 2^(5/2-3/2) =2^2/2 =2^1 =2

OpenStudy (anonymous):

@ fjugiwadi I think that @ superpik123 deserves a medal if you can spare one.

OpenStudy (anonymous):

well thanx for ur concern @robtobey